A constant force of 2.50 N accelerates a stationary particle if mass 15g through a displacement of 2.50 m . find the work done and the average power delivered.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Work done = Force x displacement.
                  = 2.50 x 2.50
                 = 6.25 Joule.
Avg. Power = Force x velocity...................(1)
From the work energy theorem,
2.50 x 2.50 = 0.5 * 15/1000 * v^2
12.5 * 1000/15 = v^2
v = root(12500/15) = > 28.86 m/s.
put v in eq. (1) we get,
Avg. Power = 2.50 x 28.86 => 721.15 Watt.